# Optimizing DINOv2 with Registers for Face Anti-Spoofing

### v1

@@ 1 Introduction
Face recognition, which identifies individuals based on the shape and position of their facial features, offers significant advantages. These include its low cost, requiring only a standard camera for image capture, and its high convenience, enabling contactless and unconstrained identification [12]. The face recognition systems have been designed to be robust against the environmental variations such as head pose, illumination, and image blur during capture, since face images can be easily changed by such variations. Consequently, a malicious actor can exploit the face recognition systems by presenting a face photo of a registered user, potentially bypassing the authentication process. As many face images of individuals are readily available online, such “spoofing attacks” have emerged as a realistic threat to face recognition systems [17].
@@
Spoofing attacks can be broadly categorized into “physical attacks” and “digital attacks”. Physical attacks include 2D attacks such as prints, replays, and cutouts, as well as 3D attacks involving masks. Digital attacks, on the other hand, encompass digital forgeries such as identity manipulation, adversarial examples, and generated images by Generative Adversarial Networks (GAN) [8]. As discussed, despite the inherent robustness of face recognition systems to environmental variations, it is crucial to detect these spoofing attacks prior to the authentication process to ensure system security.
@@
To effectively counter spoofing attacks, it is essential to detect minute differences between live and spoofed face images. It is critical to extract features from the input image that reveal global or local discrepancies, such as the texture of paper, a display device, depth, etc. Many methods [24, 23, 20, 22, 3, 21, 11, 27, 7, 9] utilizing Convolutional Neural Networks (CNNs) and Vision Transformers (ViT) [5] have been proposed to extract features inherent to spoofing attacks. Although these methods achieve high detection accuracy for simple spoofing attacks, their performance tends to degrade significantly when faced with unknown spoofing attacks not included in the training data.
@@
In this paper, we propose a novel spoofing attack detection method utilizing DINOv2 [19], enabling the capture of fine-grained details within a face image. We employ DINOv2 with registers [4] to enhance the generalizability of extracted features and to suppress perturbations in the attention mechanism, thereby enabling focused attention on essential and minute features. Furthermore, to maintain the generalization capability of the extracted features, only the last encoder block of DINOv2 is made trainable. We demonstrate the effectiveness of the proposed method through extensive experiments conducted on the dataset provided by “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025” [15], hereafter referred to as the FAS Workshop dataset, and the SiW dataset [16].
@@ 2 Related Work
This section presents a brief overview of face spoofing attack detection methods proposed with intra-dataset testing. Major methods for detecting face spoofing attacks utilize CNNs and ViT. There are several CNN-based methods such as CDCN++ [24], NAS-FAS [23], and PatchNet [20]. CDCN++ [24] extracts depth features of face images using CNN and detects spoofing attacks using these features. NAS-FAS [23] detects spoofing attacks using an optimal network derived from Neural Architecture Search (NAS). PatchNet [20] employs a patch-based approach, dividing face images into patches and extracting local features from each patch using CNN to detect spoofing attacks. CNNs cannot always recognize the various changes that occur in a face image due to spoofing attacks. In particular, the detection accuracy is degraded for unknown spoofing attacks that are not included in the training data. ViT [5] can extract both local and global features from an image, and thus can detect local and global changes in a face image caused by spoofing attacks. TransFAS [21] improved the accuracy of spoofing attack detection by adding a depth feature extraction module to ViT. Watanabe et al. [22] introduced the data augmentation method, Patch-wise Data Augmentation (PDA), which is dedicated to detecting spoofing attacks, into ViT to improve the accuracy of spoofing attack detection using ViT. Chen et al. [3] proposed a spoofing attack detection method using Segment Anything Model (SAM) [10], which is a foundation model of image segmentation based on ViT. Feng et al. [7] proposed a spoofing attack detection method that uses the features extracted from the intermediate layers of ViT and employs data augmentation methods enhancing the performance of face spoofing attack detection: Patch-wise Data Augmentation (PDA) [22] and Face Anti-Spoofing data augmentation (FAS-Aug) [2].
@@ 3.1 Overview
Fig. 1 illustrates the network architecture based on DINOv2 ViT-B/14 with registers [4]. This architecture takes an input image of $224\times 224$ pixels, divides it into $14\times 14$ patches, and feeds them into an encoder consisting of 12 layers. We use pre-trained parameters as initial values and then fine-tune the model on the training data of a face spoofing attack detection dataset. To prevent overfitting on the training data and maintain generalization performance, we unfreeze and fine-tune only the parameters of the last encoder block of DINOv2. The effectiveness of this partial unfreezing strategy for the last encoder block has been empirically confirmed, with further details provided in Sect. 4.1.4. A classification head is appended after the last layer of DINOv2 with registers. For training, we utilize a 2-class classification loss applied to the class token output from the classification head. For inference, the live probability is calculated from the output of the classification head.
@@ 3.2 DINOv2 with Registers
The proposed method employs DINOv2 with registers to effectively capture the subtle and intrinsic differences between live and spoofed images. In detecting spoofing attacks, the differences in fine textures and patterns between live and spoofed images directly affect the detection performance. Consequently, if noise appears in the attention map of ViT, it may prevent ViT from learning of essential and fine-grained features indispensable for effective spoofing attack detection. Prior research visualizing the internal operations of ViT has observed a “spike phenomenon” where abnormally high attention is focused on non-essential patches, such as background regions [4]. Such perturbations in the attention maps is attributed to outlier tokens within a portion of the embeddings output from the model, which exhibit extremely high vector norms. It has been reported that they contain global information about the entire image, but lack local features. This phenomenon is considered to be due to the reuse of unnecessary patches as temporary memory storage in ViT. To address this issue, we introduce register tokens [4] in ViT. A register is a learnable token that is treated as part of the input sequence, similar to patch and class tokens in Transformer [5]. In DINOv2 ViT-B/14 with registers, four register tokens are introduced to provide a structure dedicated to the retention and aggregation of temporary information inside the model, thereby suppressing attention disturbances caused by the reuse of unnecessary patches. Thus, by using DINOv2 with registers, the method suppresses perturbations in the attention mechanism and allows for focused attention on essential and fine-grained features.
@@ 4 Experiments and Discussion
This section describes the experiments conducted to demonstrate the effectiveness of the proposed method in detecting face spoofing attacks. The experiments use two distinct datasets to comprehensively evaluate the performance. First, we use the FAS Workshop dataset [15] provided by “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025.” Second, we use the publicly available SiW dataset [16] to allow for a comparison with existing methods. The following discussion is divided into two parts, with each section dedicated to one of these datasets.
@@ 4.1.1 Dataset
The FAS Workshop dataset consists of a variety of spoofing attacks: “Live”, “Print”, “Replay”, “Cutouts”, “Transparent”, “Plaster”, “Resin”, “Attribute-Edit”, “Face-Swap”, “Video-Driven”, “Pixel-Level”, “Semantic-Level”, “ID Consistent”, “Style Transfer”, and “Prompt Driven”. The detailed description of this dataset is available in [15]. It is an enhanced version of UniAttackData [6], which includes datasets such as [25, 26, 13, 14]. Fig. 2 shows the distribution of the classes in training dataset and validation dataset, revealing a strong class imbalance.
@@ 4.1.2 Experimental Condition
We describe the experimental conditions in the experiments. The input images are normalized with training dataset statistical characteristics, namely average and standard deviation. The image is divided into patches and input to the proposed method. As described in Sect. 3.1, the proposed method uses the pre-trained model of DINOv2 ViT-B/14 with registers [4] and performs fine tuning on the FAS Workshop dataset. The size of the mini-batch is 32, the learning rate for the binary classification head is $5\times 10^{-5}$, the learning rate for the DINOv2 backbone is $5\times 10^{-6}$, the number of epochs is 200 with early stopping with a patience of 20 epochs, and AdamW is used as optimizer. We conduct two experiments to demonstrate the effectiveness of the proposed method. The first experiment is the ablation study on unfreezing layers for DINOv2 with registers (Sect. 4.1.4). The second experiment is a comparison between the baseline method (AjianLiu) in “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025” and the proposed method (Sect. 4.1.5).
@@ 4.2.1 Dataset
SiW consists of 4,478 videos taken from 165 subjects. For each subject, there are 8 real videos and up to 20 spoofed videos. Each frame is $1,920\times 1,080$ pixels and each video is captured at 30 fps for about 15 seconds. The “Live” videos are captured under various lighting conditions while the subjects move their heads back and forward, rotate their heads, and change their facial expressions. The print attack uses two types of paper, and the display attack uses four types of display devices. For a detailed explanation of Protocols of SiW dataset, please refer to [16].
@@ 4.2.2 Data Augmentation
Since SiW only includes physical attack types, augmentation techniques can be used to great effect during training. In this experiment, we employ two data augmentation methods: a traditional method and FAS-Aug [2]. The traditional data augmentation method refers to simple, non FAS-specific alterations such as random rotation and color jitter. FAS-Aug [2] consists of 8 types of data augmentation that emulate photography noise, print attack, and display attack. As photography noise, (a) hand trembling simulation, (b) low-resolution simulation, and (c) color diversity simulation are added to the image. Since these noises occur in both live and spoofed images, the labels are not replaced when these noises are added. As print-attack noise, (d) color distortion simulation, (e) SFC-halftone artifacts, and (f) BN-halftone artifacts are added to the image. When these are added, the label is replaced by “Spoof.” As display-attack noise, (g) specular reflection artifacts and (h) moiré pattern artifacts are added to the image. When these are added, the label is replaced by “Spoof.” The parameters used in FAS-Aug are in accordance with [2]. The proposed method randomly selects one of the 8 data augmentations and applies it to the input image. Fig. 3 shows the example of images to which the 8 data augmentations included in FAS-Aug are applied.
@@ 4.2.3 Experimental Condition
The images in SiW need to be processed before conducting the experiments. Since videos in SiW [16] contain objects that can easily be recognized as spoofing attacks, face regions are extracted from each frame of the videos and used as input images. Since the bounding box for face regions is known, the face regions are extracted based on these coordinates and resized to $224\times 224$ pixels. The proposed method assumes that the input is a single image, not a video. Therefore, 5 frames are randomly extracted from the training or evaluation videos for each evaluation protocol of SiW, and used as input images for training or evaluation in the experiments as in [1]. Similarly to Sect. 4.1.2, the input images are normalized with training dataset characteristics, and the same model and parameters are used. However, we made a few adjustments due to the characteristics of SiW, which has less diverse and challenging attack types than the FAS Workshop dataset. We unfreeze the entire model for training, and switch the loss function to cross entropy and the optimizer to Nesterov Stochastic Gradient Descent [18]. For data augmentation, we systematically apply a traditional method, and use FAS-Aug with a probability of 0.2. We conduct a comparison between conventional methods and the proposed method to evaluate the effectiveness of the proposed method in detecting face spoofing attacks (Sect. 4.2.5).
@@ 4.2.4 Evaluation Metrics
1 CDCN++ [24] 0.07 0.17 0.12
NAS-FAS [23] 0.07 0.17 0.12
PatchNet [20] 0.00 0.00 0.00
Watanabe [22] 0.11 0.08 0.10
TransFAS [21] 0.00 0.00 0.00
MFAE [27] 0.00 0.00 0.00
Li [11] 0.00 0.16 0.08
Feng [7] 0.1 0.08 0.09
@@
PatchNet [20] $\pm0.00$ $\pm0.00$ $\pm0.00$
Watanabe [22] $\pm0.01$ $\pm0.01$ $\pm0.01$
TransFAS [21] $\pm0.00$ $\pm0.00$ $\pm0.00$
MFAE [27] $\pm0.00$ $\pm0.00$ $\pm0.00$

### v2

@@ 1 Introduction
Face recognition, which identifies individuals based on the shape and position of their facial features, offers significant advantages. These include its low cost, requiring only a standard camera for image capture, and its high convenience, enabling contactless and unconstrained identification [Handbook-Face-Recognition]. The face recognition systems have been designed to be robust against the environmental variations such as head pose, illumination, and image blur during capture, since face images can be easily changed by such variations. Consequently, a malicious actor can exploit the face recognition systems by presenting a face photo of a registered user, potentially bypassing the authentication process. As many face images of individuals are readily available online, such “spoofing attacks” have emerged as a realistic threat to face recognition systems [Handbook-Anti-Spoofing].
@@
Spoofing attacks can be broadly categorized into “physical attacks” and “digital attacks”. Physical attacks include 2D attacks such as prints, replays, and cutouts, as well as 3D attacks involving masks. Digital attacks, on the other hand, encompass digital forgeries such as identity manipulation, adversarial examples, and generated images by Generative Adversarial Networks (GAN) [Goodfellow-CACM-2020]. As discussed, despite the inherent robustness of face recognition systems to environmental variations, it is crucial to detect these spoofing attacks prior to the authentication process to ensure system security.
@@
To effectively counter spoofing attacks, it is essential to detect minute differences between live and spoofed face images. It is critical to extract features from the input image that reveal global or local discrepancies, such as the texture of paper, a display device, depth, etc. Many methods [Yu-CVPR-2020, Yu-PAMI-2020, Wang-CVPR-2022, Watanabe-APSIPA-2022, Chen-CoRR-2023, Wang-TBIOM-2022, Li-NN-2024, Zheng-IFS-2024, Feng-CVPRW-2025, He-CVPRW-2024] utilizing Convolutional Neural Networks (CNNs) and Vision Transformers (ViT) [Dosovitskiy-ICLR-2021] have been proposed to extract features inherent to spoofing attacks. Although these methods achieve high detection accuracy for simple spoofing attacks, their performance tends to degrade significantly when faced with unknown spoofing attacks not included in the training data.
@@
In this paper, we propose a novel spoofing attack detection method utilizing DINOv2 [Oquab-TMLR-2024], enabling the capture of fine-grained details within a face image. We employ DINOv2 with registers [Darcet-ICLR-2024] to enhance the generalizability of extracted features and to suppress perturbations in the attention mechanism, thereby enabling focused attention on essential and minute features. Furthermore, to maintain the generalization capability of the extracted features, only the last encoder block of DINOv2 is made trainable. We demonstrate the effectiveness of the proposed method through extensive experiments conducted on the dataset provided by “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025” [Liu-CoRR-2025], hereafter referred to as the FAS Workshop dataset, and the SiW dataset [Liu-CVPR-2018].
@@ 2 Related Work
This section presents a brief overview of face spoofing attack detection methods proposed with intra-dataset testing. Major methods for detecting face spoofing attacks utilize CNNs and ViT. There are several CNN-based methods such as CDCN++ [Yu-CVPR-2020], NAS-FAS [Yu-PAMI-2020], and PatchNet [Wang-CVPR-2022]. CDCN++ [Yu-CVPR-2020] extracts depth features of face images using CNN and detects spoofing attacks using these features. NAS-FAS [Yu-PAMI-2020] detects spoofing attacks using an optimal network derived from Neural Architecture Search (NAS). PatchNet [Wang-CVPR-2022] employs a patch-based approach, dividing face images into patches and extracting local features from each patch using CNN to detect spoofing attacks. CNNs cannot always recognize the various changes that occur in a face image due to spoofing attacks. In particular, the detection accuracy is degraded for unknown spoofing attacks that are not included in the training data. ViT [Dosovitskiy-ICLR-2021] can extract both local and global features from an image, and thus can detect local and global changes in a face image caused by spoofing attacks. TransFAS [Wang-TBIOM-2022] improved the accuracy of spoofing attack detection by adding a depth feature extraction module to ViT. Watanabe et al. [Watanabe-APSIPA-2022] introduced the data augmentation method, Patch-wise Data Augmentation (PDA), which is dedicated to detecting spoofing attacks, into ViT to improve the accuracy of spoofing attack detection using ViT. Chen et al. [Chen-CoRR-2023] proposed a spoofing attack detection method using Segment Anything Model (SAM) [Kirillov-ICCV-2023], which is a foundation model of image segmentation based on ViT. Feng et al. [Feng-CVPRW-2025] proposed a spoofing attack detection method that uses the features extracted from the intermediate layers of ViT and employs data augmentation methods enhancing the performance of face spoofing attack detection: Patch-wise Data Augmentation (PDA) [Watanabe-APSIPA-2022] and Face Anti-Spoofing data augmentation (FAS-Aug) [Cai-IJCV-2024].
@@ 3.1 Overview
Fig. 1 illustrates the network architecture based on DINOv2 ViT-B/14 with registers [Darcet-ICLR-2024]. This architecture takes an input image of $224\times 224$ pixels, divides it into $14\times 14$ patches, and feeds them into an encoder consisting of 12 layers. We use pre-trained parameters as initial values and then fine-tune the model on the training data of a face spoofing attack detection dataset. To prevent overfitting on the training data and maintain generalization performance, we unfreeze and fine-tune only the parameters of the last encoder block of DINOv2. The effectiveness of this partial unfreezing strategy for the last encoder block has been empirically confirmed, with further details provided in Sect. 4.1.4. A classification head is appended after the last layer of DINOv2 with registers. For training, we utilize a 2-class classification loss applied to the class token output from the classification head. For inference, the live probability is calculated from the output of the classification head.
@@ 3.2 DINOv2 with Registers
The proposed method employs DINOv2 with registers to effectively capture the subtle and intrinsic differences between live and spoofed images. In detecting spoofing attacks, the differences in fine textures and patterns between live and spoofed images directly affect the detection performance. Consequently, if noise appears in the attention map of ViT, it may prevent ViT from learning of essential and fine-grained features indispensable for effective spoofing attack detection. Prior research visualizing the internal operations of ViT has observed a “spike phenomenon” where abnormally high attention is focused on non-essential patches, such as background regions [Darcet-ICLR-2024]. Such perturbations in the attention maps is attributed to outlier tokens within a portion of the embeddings output from the model, which exhibit extremely high vector norms. It has been reported that they contain global information about the entire image, but lack local features. This phenomenon is considered to be due to the reuse of unnecessary patches as temporary memory storage in ViT. To address this issue, we introduce register tokens [Darcet-ICLR-2024] in ViT. A register is a learnable token that is treated as part of the input sequence, similar to patch and class tokens in Transformer [Dosovitskiy-ICLR-2021]. In DINOv2 ViT-B/14 with registers, four register tokens are introduced to provide a structure dedicated to the retention and aggregation of temporary information inside the model, thereby suppressing attention disturbances caused by the reuse of unnecessary patches. Thus, by using DINOv2 with registers, the method suppresses perturbations in the attention mechanism and allows for focused attention on essential and fine-grained features.
@@ 4 Experiments and Discussion
This section describes the experiments conducted to demonstrate the effectiveness of the proposed method in detecting face spoofing attacks. The experiments use two distinct datasets to comprehensively evaluate the performance. First, we use the FAS Workshop dataset [Liu-CoRR-2025] provided by “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025.” Second, we use the publicly available SiW dataset [Liu-CVPR-2018] to allow for a comparison with existing methods. The following discussion is divided into two parts, with each section dedicated to one of these datasets.
@@ 4.1.1 Dataset
The FAS Workshop dataset consists of a variety of spoofing attacks: “Live”, “Print”, “Replay”, “Cutouts”, “Transparent”, “Plaster”, “Resin”, “Attribute-Edit”, “Face-Swap”, “Video-Driven”, “Pixel-Level”, “Semantic-Level”, “ID Consistent”, “Style Transfer”, and “Prompt Driven”. The detailed description of this dataset is available in [Liu-CoRR-2025]. It is an enhanced version of UniAttackData [Fang-IJCAI-2024], which includes datasets such as [Zhang-CVPR-2019, Zhang-TBIOM-2020, Liu-WACV-2021, Liu-IFS-2022]. Fig. 2 shows the distribution of the classes in training dataset and validation dataset, revealing a strong class imbalance.
@@ 4.1.2 Experimental Condition
We describe the experimental conditions in the experiments. The input images are normalized with training dataset statistical characteristics, namely average and standard deviation. The image is divided into patches and input to the proposed method. As described in Sect. 3.1, the proposed method uses the pre-trained model of DINOv2 ViT-B/14 with registers [Darcet-ICLR-2024] and performs fine tuning on the FAS Workshop dataset. The size of the mini-batch is 32, the learning rate for the binary classification head is $5\times 10^{-5}$, the learning rate for the DINOv2 backbone is $5\times 10^{-6}$, the number of epochs is 200 with early stopping with a patience of 20 epochs, and AdamW is used as optimizer. We conduct two experiments to demonstrate the effectiveness of the proposed method. The first experiment is the ablation study on unfreezing layers for DINOv2 with registers (Sect. 4.1.4). The second experiment is a comparison between the baseline method (AjianLiu) in “The 6th Face Anti-Spoofing Workshop: Unified Physical-Digital Attacks Detection@ICCV2025” and the proposed method (Sect. 4.1.5).
@@ 4.2.1 Dataset
SiW consists of 4,478 videos taken from 165 subjects. For each subject, there are 8 real videos and up to 20 spoofed videos. Each frame is $1,920\times 1,080$ pixels and each video is captured at 30 fps for about 15 seconds. The “Live” videos are captured under various lighting conditions while the subjects move their heads back and forward, rotate their heads, and change their facial expressions. The print attack uses two types of paper, and the display attack uses four types of display devices. For a detailed explanation of Protocols of SiW dataset, please refer to [Liu-CVPR-2018].
@@ 4.2.2 Data Augmentation
Since SiW only includes physical attack types, augmentation techniques can be used to great effect during training. In this experiment, we employ two data augmentation methods: a traditional method and FAS-Aug [Cai-IJCV-2024]. The traditional data augmentation method refers to simple, non FAS-specific alterations such as random rotation and color jitter. FAS-Aug [Cai-IJCV-2024] consists of 8 types of data augmentation that emulate photography noise, print attack, and display attack. As photography noise, (a) hand trembling simulation, (b) low-resolution simulation, and (c) color diversity simulation are added to the image. Since these noises occur in both live and spoofed images, the labels are not replaced when these noises are added. As print-attack noise, (d) color distortion simulation, (e) SFC-halftone artifacts, and (f) BN-halftone artifacts are added to the image. When these are added, the label is replaced by “Spoof.” As display-attack noise, (g) specular reflection artifacts and (h) moiré pattern artifacts are added to the image. When these are added, the label is replaced by “Spoof.” The parameters used in FAS-Aug are in accordance with [Cai-IJCV-2024]. The proposed method randomly selects one of the 8 data augmentations and applies it to the input image. Fig. 3 shows the example of images to which the 8 data augmentations included in FAS-Aug are applied.
@@ 4.2.3 Experimental Condition
The images in SiW need to be processed before conducting the experiments. Since videos in SiW [Liu-CVPR-2018] contain objects that can easily be recognized as spoofing attacks, face regions are extracted from each frame of the videos and used as input images. Since the bounding box for face regions is known, the face regions are extracted based on these coordinates and resized to $224\times 224$ pixels. The proposed method assumes that the input is a single image, not a video. Therefore, 5 frames are randomly extracted from the training or evaluation videos for each evaluation protocol of SiW, and used as input images for training or evaluation in the experiments as in [Boulkenafet-FG-2017]. Similarly to Sect. 4.1.2, the input images are normalized with training dataset characteristics, and the same model and parameters are used. However, we made a few adjustments due to the characteristics of SiW, which has less diverse and challenging attack types than the FAS Workshop dataset. We unfreeze the entire model for training, and switch the loss function to cross entropy and the optimizer to Nesterov Stochastic Gradient Descent [Nesterov-USSR-1983]. For data augmentation, we systematically apply a traditional method, and use FAS-Aug with a probability of 0.2. We conduct a comparison between conventional methods and the proposed method to evaluate the effectiveness of the proposed method in detecting face spoofing attacks (Sect. 4.2.5).
@@ 4.2.4 Evaluation Metrics
1 CDCN++ [Yu-CVPR-2020] 0.07 0.17 0.12
NAS-FAS [Yu-PAMI-2020] 0.07 0.17 0.12
PatchNet [Wang-CVPR-2022] 0.00 0.00 0.00
Watanabe [Watanabe-APSIPA-2022] 0.11 0.08 0.10
TransFAS [Wang-TBIOM-2022] 0.00 0.00 0.00
MFAE [Zheng-IFS-2024] 0.00 0.00 0.00
Li [Li-NN-2024] 0.00 0.16 0.08
Feng [Feng-CVPRW-2025] 0.1 0.08 0.09
@@
PatchNet [Wang-CVPR-2022] $\pm0.00$ $\pm0.00$ $\pm0.00$
Watanabe [Watanabe-APSIPA-2022] $\pm0.01$ $\pm0.01$ $\pm0.01$
TransFAS [Wang-TBIOM-2022] $\pm0.00$ $\pm0.00$ $\pm0.00$
MFAE [Zheng-IFS-2024] $\pm0.00$ $\pm0.00$ $\pm0.00$